OpenStudy (abbycross167):

Can someone please help me answer this question? Please? I would really appreciate it... Question~ Solve the systems using the substitution method. -3x + y = 0 3x - y = -12

hartnn (hartnn):

-3x + y = 0 add 3x on both sides, what u get?

OpenStudy (abbycross167):

So it should look like this -3x+y-3x=0-3x? @hartnn I am very sorry my WiFi messed up so I didn't see your reply... I'm really sorry

hartnn (hartnn):

hey no problem :) and we had to add 3x, not subtract. -3y + y + 3y = 0 +3y so what does y= .. .

OpenStudy (abbycross167):

oh ok.. y=0?

hartnn (hartnn):

i meant, -3x + y + 3x = 0 + 3x so y= 3x now we "substitute" this in 2nd equation,

hartnn (hartnn):

3x -y = -12 substitute y= 3x in this

OpenStudy (abbycross167):

oooh ok... Sorry math is not my strong point.

hartnn (hartnn):

so what do u get?

hartnn (hartnn):